Which of the following statements are true? Please justify the answers.
1. For any two sets A and B, A∩Bᶜ=A\B.
2. The matrix [1 1 is singular.
0 0]
3. The contrapositive of '∃ y∈Z such that P(y) is true' is '∃ x∈Z such that P(x) is true'.
4. The system 2x-3y=1 and 6y-4x+2=0 has a unique solution.
5. If x, y∈C such that x²=y and y²=x, then x=y=1.
1
Expert's answer
2019-07-09T12:08:38-0400

1. Relationship between relative and absolute complements:


"A\setminus B=A\cap B^C"

Therefore, the first statement is true.


2. A matrix is singular iff its determinant is 0.


"\begin{vmatrix}
 1 & 1 \\
 0 & 0
\end{vmatrix}=0-0=0"

Therefore, the second statement is true.


3. A conditional statement has a contrapositive, a converse, and an inverse. The satement  '∃ y∈Z such that P(y) is true' is an existential statement and may have only negation.


Therefore, the third statement is not true.


4. Rewrite the system


"2x-3y=1""4x-6y=2"

The system has infinitely many solutions:


"x={3\over 2}y+{1 \over 2}, y\in\R"

Therefore, the fourth statement is not true.


5. Rewrite the system


"x^2=y""y^2=x"

Then


"x=y^2 =x^4""x^4-x=0""x(x-1)(x^2+x+1)=0"

"x_1=0, x_2=1, x_3={-1\over 2}-{i\sqrt 3 \over 2}, x_4={-1\over 2}+{i\sqrt 3 \over 2}"

"y_1=0, y_2=1, y_3={-1\over 2}+{i\sqrt 3 \over 2}, y_4={-1\over 2}-{i\sqrt 3 \over 2}"

Therefore, the fifth statement is not true.